# PVF (Parameter Vulnerability Factor): A Quantitative Metric Measuring AI Vulnerability and Resilience Against Parameter Corruptions

### v1

@@ Abstract
Reliability of AI systems is a fundamental concern for the successful deployment and widespread adoption of AI technologies. Unfortunately, the escalating complexity and heterogeneity of AI hardware systems make them inevitably and increasingly susceptible to hardware faults (e.g., bit flips) that can potentially corrupt model parameters. Given this challenge, this paper aims to answer a critical question: How likely is a parameter corruption to result in an incorrect model output? To systematically answer this question, we propose a novel quantitative metric, Parameter Vulnerability Factor (PVF), inspired by architectural vulnerability factor (AVF) in computer architecture community, aiming to standardize the quantification of AI model resilience/vulnerability against parameter corruptions. We define a model parameter’s PVF as the probability that a corruption in that particular model parameter will result in an incorrect output. Similar to AVF, this statistical concept can be derived from statistically extensive and meaningful fault injection (FI) experiments. In this paper, we present several use cases on applying PVF to three types of tasks/models during inference – recommendation (DLRM), vision classification (CNN), and text classification (BERT), while presenting an in-depth resilience analysis on DLRM workload as it occupies the most inference cycles in Meta’s data center. In DLRM, our FI results show that top-MLP layers in DLRM are the most vulnerable parameter component, while embedding tables exhibit remarkable resilience to parameter corruptions. We also discuss the potential use case of PVF during model training. PVF can provide pivotal insights to AI hardware designers in balancing the tradeoff between fault protection and performance/efficiency such as mapping vulnerable AI parameter components to well-protected hardware modules. PVF metric is applicable to any AI model and has a potential to help unify and standardize AI vulnerability/resilience evaluation practice.
@@ I Introduction
  • Parameter-level Quantitative Assessment: As a quantitative metric, PVF focuses on parameter-level vulnerability, measuring the probability that a corruption in a given model parameter will result in an incorrect model output. The PVF concept draws inspiration from the Architectural Vulnerability Factor (AVF) in the computer architecture community, where an architectural structure’s AVF is the probability that a fault in that particular structure will result in a program output error.
@@
  • Standardization of AI Resilience Evaluation: We propose to use PVF as a standard metric for AI resilience/vulnerability evaluation. It has the potential to unify and standardize such practices, making it easier to compare the reliability of different AI systems/parameters and fostering open collaboration and progress in the industry/research community.
@@ III-B Fault Model
In this paper, we consider “a parameter corruption” using three fault models. These simulate hardware faults such as memory bit flips due to soft errors [7]. These fault models are used in previous studies [21, 17, 13]. We acknowledge different hardware platforms have different error patterns, and the fault model here may not fully represent a realistic hardware error pattern. Nevertheless, PVF is applicable to any given fault model.
@@ IV-C PVF under MBF
Observation 1: DLRM exhibits strong resilience against parameter corruption. We can see that when there is 1 bit flip, all the parameter components have low PVF ($<0.5\%$). That means for every 1000 FI experiments, fewer than 5 output will be incorrect. Alternatively stated, even if a parameter corruption occurs, the probability that an incorrect output is observed is less than 0.5%. which represents a strong resilience of DLRM to parameter corruption. Just like other neural network models such as MLP and CNNs [12, 16], our results show that DLRM possess strong resilience. This is because certain neural network components can mask certain set of errors due to their intrinsic operations. For example, a non-linear activation function like ReLu will naturally map any negative input to 0 regardless of what the negative input is. Another reason is that neural network model itself is probabilistic, e.g., if the softmax probability output is modified by a small value, the resulting inference classification may be correct.
Observation 2: DLRM embedding tables show strong resilience against parameter corruption. We observe that even with 128 bit flips during each inference, the PVF of embedding table is still close to 0 ($<0.0001\%$), exhibiting remarkable resilience against corruptions; for MLP components, PVF has increased to 40% or 10% for top-MLP and bot-MLP components respectively, while observing multiple NaN values. This is attributed to embedding tables being highly sparse, and parameter corruptions are only activated when the particular corrupted parameter is “hit” by the corresponding sparse feature. Even if activated, it may get masked by the subsequent neural processing. Consequently, a bit flip in the embedding table is likely to not get exposed or reflected in the DLRM output in most cases. This further lowers the probability of error exposure for the architecture under study. Note that, because different parameter components has different parameter count, the same bit flip count is essentially equivalent to different BERs for these parameter components. For example, 1000 bit flips for embedding table is 0.000185%, and for top-MLP and bot-MLP are 0.641% and 0.3124%. Thus, when we compare the PVF across different parameter components, we need to explicitly state the underlying fault model (which can be specified by users based on the assumption regarding hardware faults) to make a fair and clear comparison.
IV-D PVF under MBBF
@@
Observation 3: DLRM embedding tables show strong resilience under burst errors. Fig. 4 presents the PVF under burst errors, where we sweep the burst error counts from 1 to 128. In this case, each burst error will flip two consecutive bits. Results show that, even at 128 burst errors, the PVF of embedding table is still close to 0 ($<0.0001\%$), while for top mlp and bottom mlp, the PVF increased to 50% and 30% respectively under 128 burst errors, higher than the PVF under MBF which is intuitively reasonable.
IV-E PVF under SBF
@@
Observation 4: DLRM MLP layers have higher vulnerability (PVF) than embedding tables. Based on Fig. 5, we can also see that MLP components have higher PVF than embedding tables. The top-MLP component has highest PVF among all target parameter components, while bot-MLP component has higher PVF than most embedding tables. The reason for this is that embedding tables are highly sparse (stated in Observation 2 and 3 earlier). Consequently, a bit flip in the embedding table is likely not to get exposed or reflected in the DLRM processing in most cases. This further lowers the probability of error reflection.
Observation 5: Top-MLP are more vulnerable than bot-MLP. We also observe that top MLP component has higher PVF than bottom MLP. This is attributed to the top-MLP being closer to the final model, and hence has less of a chance to be mitigated by inherent error masking probability of DLRM models. Note that in this case, top-MLP has more parameters than bot-MLP – meaning that same bit flip count suggests smaller BER for top-MLP.
Observation 6: Under SBF fault model, the PVF of individual embedding table is highly (inversely) correlated with the parameter counts. Under SBF fault model, we observe that the PVF of embedding tables varies by orders of magnitude, and is highly (inversely) correlated with the parameter count (as seen in Fig. 6), which shows the inverse of the parameter count. We draw a correlation that, the smaller an embedding table is, the higher its PVF (under SBF model). This is intuitively reasonable because under SBF fault model, a single bit flip will always occur in the target parameter regardless of the size of the target parameter. Thus, for an embedding table with smaller parameter count, it is more likely the bit flip will be activated by the sparse feature. However, it is worth to note that, for a smaller embedding table, the probability of a bit flip is also likely to be less than a larger embedding table.
@@
Observation 7: PVF can identify vulnerable bits; we observe bit 30 is the most vulnerable bit (in FP32 DLRM). Fig. 7 presents PVF of different bit positions (we omit some least significant bits because they consistently have 0 PVF) for embedding, top-MLP, and bot-MLP. In this scenario, to obtain the PVF of a given bit position, we only inject bit flips to that specific bit position. A key observation is that the sign bit (bit 31) is not the most vulnerable bit. Bit 30 is the most vulnerable bit because compared to other bits, it is more likely to result in large values as well as abnormal data such as NaNs; in FP32 data, NaN could be caused by a “0” to “1” flip at bit position 30.
@@ VII Discussion
The PVF, which quantifies the likelihood that a corruption in a specific model parameter will result in an incorrect output, could be extended to different AI models/tasks. Note that the definition of “incorrect output” will be different for different models/tasks and can be customized to users’ need. For example, it can be applied to a large language model (LLM) for question-answering tasks, where in this specific case, an “incorrect output” is the wrong answer. It can also be customized to users’ specific needs – for example, for an GenAI model/task, users may have a specific metric to define “incorrectness” for the generated content. We leave this for future follow-up research.
PVF can also be extended to training phase to assess the impact of parameter corruptions on the model’s ability to converge effectively. During training, the model’s parameters are iteratively updated to minimize a loss function. If a parameter is corrupted, it could potentially disrupt this learning process and prevent the model from converging to an optimal solution. By applying the PVF concept during training, we could quantify the likelihood that a corruption in each parameter would lead to such a failure to converge. One important difference between training and inference is that during training, the parameters themselves are under a dynamic update and hence the PVF is potentially a function of timing as well. This will need further investigation and experimentation.
Nevertheless, applying and obtaining PVF during training can provide valuable insights into the robustness of the training process. For instance, if certain parameters have a high PVF, it would suggest that corruptions in these parameters are particularly detrimental to the model’s learning. This information could guide efforts to improve the fault tolerance of the training process, such as by developing training algorithms that can recover from parameter corruptions, or by implementing hardware or software protections for the most vulnerable parameters.
Moreover, PVF could be used to assess the trade-off between training robustness and computational efficiency. Training AI models, particularly large models like BERT or DLRM, can be computationally intensive and time-consuming. Understanding the PVF of the model’s parameters could help balance the need for robustness against hardware faults with the desire to minimize computational resources and training time.
@@ VIII Conclusion
In this study, we introduce PVF, a metric that measures the likelihood of a model parameter corruption leading to an erroneous output. To obtain this metric, we conduct comprehensive fault injection (FI) experiments on three different tasks/models during inference – recommendation (DLRM), vision classification (CNN), and text classification (BERT). The PVF outcomes allow us to examine the resilience and vulnerabilities of various parameter components within a specific model. The introduction of the PVF metric offers crucial insights for AI hardware designers, aiding them in striking a balance between fault protection and performance/efficiency, such as assigning vulnerable AI parameter components to highly protected hardware modules. The PVF metric is versatile and can be applied to any AI model, potentially serving as a unifying and standardizing tool for evaluating AI vulnerability and resilience.

### v2

# PVF (Parameter Vulnerability Factor): A Quantitative Metric Measuring AI Vulnerability Against Parameter Corruptions

@@ Abstract
Reliability of AI systems is a fundamental concern for the successful deployment and widespread adoption of AI technologies. Unfortunately, the escalating complexity and heterogeneity of AI hardware systems make them increasingly susceptible to hardware faults (e.g., bit flips) that can potentially corrupt model parameters. When this occurs during AI inference/servicing, it can potentially lead to incorrect or degraded model output for users, ultimately affecting the quality and reliability of AI services. In light of the escalating threat, it is crucial to address key questions: How vulnerable are AI models to parameter corruptions, and how do different components (such as modules, layers) of the models exhibit varying vulnerabilities to parameter corruptions? To systematically address this question, we propose a novel quantitative metric, Parameter Vulnerability Factor (PVF), inspired by architectural vulnerability factor (AVF) in computer architecture community, aiming to standardize the quantification of AI model vulnerability against parameter corruptions. We define a model parameter’s PVF as the probability that a corruption in that particular model parameter will result in an incorrect output. Similar to AVF, this statistical concept can be derived from statistically extensive and meaningful fault injection (FI) experiments. In this paper, we present several use cases on applying PVF to three types of tasks/models during inference – recommendation (DLRM), vision classification (CNN), and text classification (BERT), while presenting an in-depth vulnerability analysis on DLRM. In DLRM, our FI results show that different parts of DLRM present different vulnerability levels: top-MLP layers are the most vulnerable parameter component, while embedding tables exhibit comparatively stronger resilience. We also discuss the potential use case of PVF during model training. PVF can provide pivotal insights to AI hardware designers in balancing the tradeoff between fault protection and performance/efficiency such as mapping vulnerable AI parameter components to well-protected hardware modules. PVF metric is applicable to any AI model and has a potential to help unify and standardize AI vulnerability/resilience evaluation practice.
@@ I Introduction
  • Parameter-level Quantitative Assessment: As a quantitative metric, PVF concentrates on parameter-level vulnerability, calculating the likelihood that a corruption in a specific model parameter will lead to an incorrect model output. This “parameter” can be defined at different scales and granularities, such as an individual parameter or a group of parameters. By applying PVF to several example models, we quantify the varying vulnerability levels of different parts of a given model.
@@
  • Standardization of AI Vulnerability/Resilience Evaluation: We propose to use PVF as a standard metric for AI vulnerability/resilience evaluation. It has the potential to unify and standardize such practices, making it easier to compare the reliability of different AI systems/parameters and fostering open collaboration and progress in the industry/research community.
@@ III-B Fault Model
In this paper, we consider parameter corruptions using three hardware fault models. These simulate hardware faults such as memory bit flips due to soft errors [7]. These fault models are used in previous studies [20, 16, 12]. We recognize that different hardware platforms exhibit different error patterns, and the fault model used here may not fully represent realistic hardware error patterns. However, PVF is adaptable to any user-specified fault model based on users’ specific scenarios.
@@ IV-C PVF under MBF
Observation 1: Different parts of DLRM present different vulnerability levels. We observe that different parts of DLRM present different vulnerability level, e.g., under a single bit flip, most embedding tables have low PVF, e.g., less than 0.0001%; however, top-MLP can have 0.4% under even a single bit flip. This is significant – for every 1000 inferences, 4 inferences will be incorrect. This highlights the importance of protecting specific vulnerable parameters for a given model based on the PVF measurement.
We observe that even with 128 bit flips during each inference, the PVF of embedding table is still low ($<0.0001\%$), exhibiting notable resilience against corruptions; for MLP components, PVF has increased to 40% or 10% for top-MLP and bot-MLP components respectively, while observing multiple NaN values. This is attributed to embedding tables being highly sparse, and parameter corruptions are only activated when the particular corrupted parameter is “hit” by the corresponding sparse feature. Even if activated, it may get masked by the subsequent neural processing. Consequently, a bit flip in the embedding table is likely to not get exposed or reflected in the DLRM output in most cases. This further lowers the probability of error exposure for the architecture under study. Note that, because different parameter components has different parameter count, the same bit flip count is essentially equivalent to different BERs for these parameter components. For example, 1000 bit flips for embedding table is 0.000185%, and for top-MLP and bot-MLP are 0.641% and 0.3124%. Thus, when we compare the PVF across different parameter components, we need to explicitly state the underlying fault model (which can be specified by users based on the assumption regarding hardware faults) to make a fair and clear comparison.
@@
Fig. 4 presents the PVF under burst errors, where we sweep the burst error counts from 1 to 128. In this case, each burst error will flip two consecutive bits. Results show that, even at 128 burst errors, the PVF of embedding table is still low ($<0.0001\%$), while for top mlp and bottom mlp, the PVF increased to 50% and 30% respectively under 128 burst errors, higher than the PVF under MBF which is intuitively reasonable.
IV-D PVF under SBF
@@
Observation 2: Top-MLP has higher vulnerability than bot-MLP and embedding tables, while different individual embedding tables show different vulnerabilities. Based on Fig. 5, we observe that MLP components have higher PVF than embedding tables. The top-MLP component has highest PVF among all target parameter components, while bot-MLP component has higher PVF than most embedding tables. The reason for this is that embedding tables are highly sparse (stated in Observation 2 and 3 earlier). Consequently, a bit flip in the embedding table is likely not to get exposed or reflected in the DLRM processing in most cases. This further lowers the probability of error reflection.
We also observe that top MLP component has higher PVF than bottom MLP. This is attributed to the top-MLP being closer to the final model, and hence has less of a chance to be mitigated by inherent error masking probability of DLRM models. Note that in this case, top-MLP has more parameters than bot-MLP – meaning that same bit flip count suggests smaller BER for top-MLP.
Under SBF fault model, we observe that the PVF of embedding tables varies by orders of magnitude, and is highly (inversely) correlated with the parameter count (as seen in Fig. 6), which shows the inverse of the parameter count. We draw a correlation that, the smaller an embedding table is, the higher its PVF (under SBF model). This is intuitively reasonable because under SBF fault model, a single bit flip will always occur in the target parameter regardless of the size of the target parameter. Thus, for an embedding table with smaller parameter count, it is more likely that the bit flip will be activated by the sparse feature. However, it is worth to note that, for a smaller embedding table, the probability of a bit flip occurrence is also smaller than a larger embedding table.
@@
Observation 3: Different bits have different vulnerability levels. Fig. 7 presents PVF of different bit positions (we omit some least significant bits because they consistently have 0 PVF) for embedding, top-MLP, and bot-MLP. In this scenario, to obtain the PVF of a given bit position, we only inject bit flips to that specific bit position. A key observation is that the sign bit (bit 31) is not the most vulnerable bit. Bit 30 is the most vulnerable bit because compared to other bits, it is more likely to result in large values as well as abnormal data such as NaNs; in FP32 data, NaN could be caused by a “0” to “1” flip at bit position 30.
@@ VII Discussion
PVF is a versatile metric that can be tailored to various AI models/tasks. The definition of an “incorrect output” will vary based on the model/task and can be adapted to suit user requirements. For instance, in a large language model (LLM) used for question-answering tasks, an incorrect output would be an incorrect answer. For a GenAI model/task, users might have a specific metric to define “incorrectness” for the generated content. We anticipate that the introduction of PVF will stimulate diverse use cases in both research and production settings.
PVF is also adaptable to various hardware fault models. While our study presents PVF under three fault models, we acknowledge that different hardware platforms under varying physical conditions may exhibit different fault models, such as bit flips, statistical variations, etc. The principle of PVF remains applicable, and the method to calculate PVF remains consistent. The only modification required is the manner in which the fault is injected, based on the assumed fault models.
Furthermore, PVF can be extended to the training phase to evaluate the effects of parameter corruptions on the model’s convergence capability. During training, the model’s parameters are iteratively updated to minimize a loss function. A corruption in a parameter could potentially disrupt this learning process, preventing the model from converging to an optimal solution. By applying the PVF concept during training, we could quantify the probability that a corruption in each parameter would result in such a convergence failure. A key distinction between training and inference is that during training, the parameters are dynamically updated, making the PVF potentially a function of time as well. We reserve this aspect for future exploration and investigation.
@@ VIII Conclusion
In this research, we propose a novel quantitative metric, the Parameter Vulnerability Factor (PVF), designed to quantify the vulnerability of AI models to parameter corruptions. Through fault injection, PVF can be calculated for any target parameter component of a given AI model. In this paper, we present three case studies where we apply PVF to recommendation (DLRM), vision classification (CNN), and text classification (BERT). The PVF measurements allow us to examine the vulnerability levels of different parameter components within a specific model. The introduction of the PVF metric provides crucial insights for AI hardware designers, assisting them in balancing fault protection with performance and efficiency. For instance, it can guide the assignment of vulnerable AI parameter components to highly protected hardware modules. The PVF metric is versatile and can be applied to any AI model. It has the potential to serve as a unifying and standardizing tool for evaluating AI vulnerability and resilience.